Doubly ionized lithium Li2+ (Z = 3) and triply ionized beryllium Be3+ (Z = 4) each emit a line spectrum. For a certain series of lines in the lithium spectrum, the shortest wavelength is 253.5 nm. For the same series of lines in the beryllium spectrum, what is the shortest wavelength?
Answer #1

Given :- 4 2t (2-3) lonized Be Be 3 (-4) - 263-50m. for bettiny. 1/2 = 2²R (/n; - 1/0) % = 2 (1091343x10+) 762% (2536x10- 9)